# Incompressible Limit of Compressible Ideal MHD Flows inside a Perfectly Conducting Wall

### v4

@@ 1.4 Organization of the paper
This paper is organized as follows. In section 2, we discuss the main difficulties and briefly introduce our strategies to tackle the problem. Then section 3 is devoted to the proof of uniform estimates in Mach number. Combined with the compactness argument, we conclude the incompressible limit in section 4.
@@ Acknowledgment.
The research of Jiawei Wang is supported by the NSFC (Grants 12071044 and 12131007). Jiawei Wang would like to thank his Ph.D. advisor Prof. Qiangchang Ju for helpful suggestions. Junyan Zhang would like to thank Chenyun Luo for the kind hospitality during his visit at The Chinese University of Hong Kong.
@@
  • (Tangential components) $\bar{u}:=(u_{1},\cdots,u_{d-1})$ and $\bar{B}:=(B_{1},\cdots,B_{d-1})$. We also write ¯:=(¯,$\overline{\nabla}:=(\overline{\partial}{}_{1},\overline{\partial}{}_{2})$ as the tangential components of the operator $\nabla$.
@@
  • (Anisotropic Sobolev norms) $*\|\cdot\|_{m,*}$: For any function $f(t,x)\text{ on }[0,T]\times\Omega$, $\|f\|_{m,*}^{2}:=\sum_{\langle\alpha\rangle\leq m}\|\partial_{*}^{\alpha}f(t,\cdot)\|_{0}^{2}$ denotes the $m$-th order space-time anisotropic Sobolev norm of $f$.
@@ 2.2 Key observation: hidden structure of Lorentz force in the vorticity analysis
First, the entropy is easy to control thanks to $D_{t}S=0$, so it suffices to analyze the relations between $(u,B)$ and $p$. We start with the control of $\|(u,B)\|_{4}$. Using div-curl decomposition, the $H^{4}$ Sobolev norms are bounded by ×$\|\nabla\times u,\nabla\times B\|_{3}$, $\|\nabla\cdot u,\nabla\cdot B\|_{3}$ and the normal traces $|u_{3},B_{3}|_{3.5}$. The boundary conditions (1.4) eliminate the normal traces, and the divergence part is reduced to tangential derivatives $\|\varepsilon^{2}D_{t}p\|_{3}$ thanks to the continuity equation and the divergence constraint $\nabla\cdot B=0$. As for the vorticity, taking curl in the momentum equation yields the evolution equation of vorticity
@@ 2.3 Reduction of pressure and design of energy functional
[…]
@@
[…]
@@ 3.1 $L^{2}$ estimate
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\varepsilon^{2}p^{2}+\rho|u|^{2}+|B|^{2}+S^{2}\,\mathrm{d}x$
$=$ $\frac{1}{2}\int_{\Omega}\nabla\cdot u(\varepsilon^{2}p^{2}+|B|^{2}+S^{2})\,\mathrm{d}x-\int_{\Omega}\nabla\cdot(pu)\,\mathrm{d}x$
$+\int_{\Omega}(B\cdot\nabla B)\cdot u+(B\cdot\nabla u)\cdot B\,\mathrm{d}x-\int_{\Omega}\frac{1}{2}\nabla|B|^{2}\cdot u+(B\cdot B)\nabla\cdot u\,\mathrm{d}x$
$=$ $\frac{1}{2}\int_{\Omega}\nabla\cdot u(\varepsilon^{2}p^{2}+|B|^{2}+S^{2})\,\mathrm{d}x,$ (3.1)
which give the energy estimate:
$\frac{\mathrm{d}}{\mathrm{d}t}\|(\varepsilon p,u,B,S)\|_{0}^{2}\leq P(E_{4}(t)).$ (3.2)
@@ 3.2.1 Reduction of pressure
Recall that $D_{t}=\partial_{t}+\bar{u}\cdot\overline{\nabla}+u_{3}\partial_{3}$ and $B\cdot\nabla=\bar{B}\cdot\overline{\nabla}+B_{3}\partial_{3}$ are both tangential derivatives thanks to $u_{3}=B_{3}=0$ on $\Sigma$. The above inequalities (3.4)-(3.5) show that $\nabla P$ is reduced to $\partial_{t}u$, ¯$\overline{\partial}{}u$, $\omega(x_{3})\partial_{3}u$, ¯$\overline{\partial}{}B$ and $\omega(x_{3})\partial_{3}B$ for some weight function $\omega(x_{3})$ vanishing on $\Sigma$. Thus, the momentum equation reduces a normal derivative of $P$ to a tangential derivative of $u,B$. Also, one can reduce the Sobolev norm of $\nabla P$ in the same way. Let $\mathcal{T}=\partial_{t}$ or $\overline{\partial}{}$ or $\omega(x_{3})\partial_{3}$ and $D=\partial$ or $\partial_{t}$. The above estimate yields the control of $\|\mathcal{T}^{\alpha}D^{k}\nabla p\|_{0}$ after taking $\mathcal{T}^{\alpha}D^{k}$, $k+\langle\alpha\rangle\geq 1$, to (3.3). Specifically, at the leading order, $\|\mathcal{T}^{\alpha}D^{k}\nabla(p+|B|^{2}/2)\|_{0}$ is controlled by
@@ 3.2.2 Div-Curl analysis
$\|X\|_{s}^{2}\lesssim\|X\|_{0}^{2}+\|\nabla\cdot X\|_{s-1}^{2}+\|\nabla\times X\|_{s-1}^{2}+|\overline{\partial}{}X\cdot N|^{2}_{s-\frac{3}{2}}.$ (3.7)
@@
The $L^{2}$ norm of $u$ and $B$ has been controlled in (3.2), while the control of $\|\varepsilon^{(k-1)_{+}}\partial_{t}^{k}(u,B)\|_{0}~{}(1\leq k\leq 3)$ is parallel to the case of $k=0$ and will be postponed to later sections. The divergence part is reduced to the estimates of $p$ by using the continuity equation
$\|\varepsilon^{(k-1)_{+}}\nabla\cdot\partial_{t}^{k}u\|_{3-k}^{2}=\left\|\varepsilon^{(k-1)_{+}+2}\partial_{t}^{k}(D_{t}p)\right\|_{3-k}^{2},$ (3.10)
@@
$\sum_{k=0}^{3}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}u\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}B\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}+1}B\times(\nabla\times\partial_{t}^{k}B)\right\|_{3-k}^{2}\right)\leq P(E_{4}(t),E_{5}(t)).$ (3.11)
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho\left|\partial^{3}\nabla\times u\right|^{2}+\left|\partial^{3}\nabla\times B\right|^{2}+\left|\varepsilon B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t),E_{5}(t)).$
@@
$\underbrace{+\int_{\Omega}(\partial^{3}\nabla\times B)\cdot\partial^{3}\nabla\times(\varepsilon^{2}BD_{t}p)\,\mathrm{d}x}_{:=\mathcal{K}}.$ (3.16)
Based on the concrete form of the commutators, a straightforward product estimate for $\mathcal{I}_{j}$ $(1\leq j\leq 4)$ gives us
@@
The most difficult term is $\mathcal{K}$ because the highest-order term has 5-th order derivative and thus cannot be controlled by $E_{4}(t)$:
$\varepsilon^{2}\partial^{3}\nabla\times(BD_{t}p)=$ $~{}\varepsilon^{2}\partial^{3}((\nabla D_{t}p)\times B+D_{t}p(\nabla\times B))$
$=$ $-\varepsilon^{2}B\times(\partial^{3}\nabla D_{t}p)\underbrace{-\varepsilon^{2}[\partial^{3},B\times](\nabla D_{t}p)+\partial^{3}(\varepsilon^{2}D_{t}p(\nabla\times B))}_{\mathcal{R}_{2}},$
where straightforward calculation shows that $\|\mathcal{R}_{2}\|_{0}\leq P(E_{4}(t))$. We then further analyze the problematic term $-\varepsilon^{2}B\times(\partial^{3}\nabla D_{t}p)$. We invoke a simple vector identity (in 3D) $(B\cdot\nabla)B-\frac{1}{2}\nabla|B|^{2}=-B\times(\nabla\times B)$ to rewrite the momentum equation to be
@@
$-\varepsilon^{2}B\times(\partial^{3}\nabla D_{t}p)=-\varepsilon^{2}B\times(\partial^{3}D_{t}\nabla p)-\varepsilon^{2}B\times(\partial^{3}(\nabla u_{j}\partial_{j}p))$
$=$ $~{}\varepsilon^{2}B\times\partial^{3}D_{t}(\rho D_{t}u)+\varepsilon^{2}B\times\partial^{3}D_{t}(B\times(\nabla\times B))-\varepsilon^{2}B\times(\partial^{3}(\nabla u_{j}\partial_{j}p))$
$=$ $~{}\varepsilon^{2}\rho B\times(\partial^{3}D_{t}^{2}u)+\underbrace{\varepsilon^{2}B\times D_{t}(B\times\partial^{3}(\nabla\times B))}_{=\mathcal{K}_{1}}$
$\underbrace{+\varepsilon^{2}B\times([\partial^{3}D_{t},\rho]D_{t}u)+\varepsilon^{2}B\times\left([\partial^{3},D_{t}](\nabla\times B)+D_{t}\left([\partial^{3},B\times](\nabla\times B)\right)\right)-\varepsilon^{2}B\times(\partial^{3}(\nabla u_{j}\partial_{j}p))}_{\mathcal{R}_{3}},$ (3.18)
@@
$\int_{\Omega}\varepsilon^{2}(\partial^{3}\nabla\times B)\cdot\left(B\times D_{t}\left(B\times\partial^{3}(\nabla\times B)\right)\right)\,\mathrm{d}x.$ (3.19)
@@
$\int_{\Omega}\varepsilon^{2}(\partial^{3}\nabla\times B)\cdot\left(B\times D_{t}\left(B\times\partial^{3}(\nabla\times B)\right)\right)\,\mathrm{d}x=-\int_{\Omega}\varepsilon^{2}D_{t}\left(B\times(\partial^{3}\nabla\times B)\right)\cdot\left(B\times(\partial^{3}\nabla\times B)\right)\,\mathrm{d}x,$ (3.20)
@@
$-\int_{\Omega}\varepsilon^{2}D_{t}\left(B\times(\partial^{3}\nabla\times B)\right)\cdot\left(B\times(\partial^{3}\nabla\times B)\right)\,\mathrm{d}x$
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\varepsilon^{2}\left|B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x+\underbrace{\frac{1}{2}\int_{\Omega}(\nabla\cdot u)\left|B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x}_{\mathcal{R}_{4}}.$ (3.21)
@@
$\mathcal{K}+\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\varepsilon^{2}\left|B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x\leq\|B\|_{4}\left(P(E_{4}(t))+\|\rho B\|_{\infty}\|\varepsilon^{2}D_{t}^{2}u\|_{3}\right)\leq P(E_{4}(t),E_{5}(t)),$ (3.22)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\|\nabla\times u\|_{3}^{2}+\|\nabla\times B\|_{3}^{2}+\left\|\varepsilon B\times(\nabla\times B)\right\|_{3}^{2}\right)\leq P(E_{4}(t),E_{5}(t)).$ (3.23)
@@
Similarly, we can prove the same conclusion for $\partial^{\alpha}\partial_{t}^{k}$ with $k+|\alpha|=3$ by replacing $\partial^{3}$ with $\varepsilon^{(k-1)_{+}}\partial^{\alpha}\partial_{t}^{k}$. Indeed, the highest order derivatives in the above commutators do not exceed 4-th order, and there is no loss of Mach number weight because none of the above steps creates negative power of Mach number. Besides, the problematic term can still be analyzed in the same manner:
$\varepsilon^{(k-1)_{+}+2}B\times(\partial^{\alpha}\partial_{t}^{k}\nabla D_{t}p)\stackrel{{\scriptstyle L}}{{=}}-\varepsilon^{(k-1)_{+}}\rho B\times\partial^{\alpha}\partial_{t}^{k}(\varepsilon^{2}D_{t}^{2}u)-\varepsilon^{(k-1)_{+}}\varepsilon^{2}B\times D_{t}\left(B\times(\partial^{\alpha}\partial_{t}^{k}(\nabla\times B))\right),$
where the $L^{2}$ norms of the omitted lower-order terms are controlled by $P(E_{4}(t))$. The contribution of the last term in the above equality is again the energy term $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\varepsilon^{2}\left|B\times(\partial^{\alpha}\partial_{t}^{k}\nabla\times B)\right|^{2}\,\mathrm{d}x$. The proof of this follows in the same way if we replace $\partial^{3}$ above by $\varepsilon^{(k-1)_{+}}\partial^{\alpha}\partial_{t}^{k}$ (that is, using the vector identity $(\mathbf{u}\times\mathbf{v})\cdot\mathbf{w}=-(\mathbf{u}\times\mathbf{w})\cdot\mathbf{v}$. Hence, we can conclude the vorticity analysis for $E_{4}(t)$ with the following inequality
$\sum_{k=0}^{3}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}u\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}B\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}+1}B\times\left(\partial_{t}^{k}\nabla\times B\right)\right\|_{3-k}^{2}\right)\leq P(E_{4}(t),E_{5}(t)).$ (3.24)
@@
$=$ $~{}\left\|\varepsilon^{4+(k-1)_{+}}\partial_{t}^{k}\mathcal{T}^{2}(D_{t}p)\right\|_{2-k}^{2}+\left\|\varepsilon^{2+(k-1)_{+}}[\mathcal{T}^{2},\nabla\cdot]\partial_{t}^{k}u\right\|_{2-k}^{2},$ (3.27)

### v6

@@ 1.4 Organization of the paper
This paper is organized as follows. In section 2, we discuss the main difficulties and briefly introduce our strategies to tackle the problem. Then section 3 is devoted to the proof of uniform estimates in Mach number. Combined with the compactness argument, we conclude the incompressible limit in section 4. In Appendix A, we prove the local existence of (1.13) for initial data in $H_{*}^{8}(\Omega)$.
@@ Acknowledgment.
The research of Jiawei Wang is supported by the National Natural Science Foundation of China (Grants 12131007 and 12071044) and the Basic Science Center Program (No: 12288201) of the National Natural Science Foundation of China. Jiawei Wang would like to thank his Ph.D. advisor Prof. Qiangchang Ju for helpful suggestions. Junyan Zhang would like to thank Chenyun Luo for the kind hospitality during his visit at The Chinese University of Hong Kong.
@@
  • (Tangential components) $\bar{u}:=(u_{1},\cdots,u_{d-1})^{\mathrm{T}}$ and $\bar{B}:=(B_{1},\cdots,B_{d-1})^{\mathrm{T}}$. We also write ¯:=(¯,$\overline{\nabla}:=(\overline{\partial}{}_{1},\cdots,\overline{\partial}{}_{d-1})^{\mathrm{T}}$ as the tangential components of the operator $\nabla$.
@@
  • (Anisotropic Sobolev norms) $\|\cdot\|_{m,*}$: For any function $f(t,x)\text{ on }[0,T]\times\Omega$, $\|f\|_{m,*}^{2}:=\sum_{\langle\alpha\rangle\leq m}\|\partial_{*}^{\alpha}f(t,\cdot)\|_{0}^{2}$ denotes the $m$-th order space-time anisotropic Sobolev norm of $f$.
@@ 2.2 Key observation: hidden structure of Lorentz force in the vorticity analysis
First, the entropy is easy to control thanks to $D_{t}S=0$, so it suffices to analyze the relations between $(u,B)$ and $p$. We start with the control of $\|(u,B)\|_{4}$. Using div-curl decomposition, the $H^{4}$ Sobolev norms are bounded by ×$\|\nabla\times u,\nabla\times B\|_{3}$, $\|\nabla\cdot u,\nabla\cdot B\|_{3}$ and the normal traces $|u_{3},B_{3}|_{3.5}$. The boundary conditions (1.4) eliminate the normal traces, and the divergence part is reduced to tangential derivatives $\|\varepsilon^{2}D_{t}p\|_{3}$ (here we temporarily abuse the notations to write $\mathcal{F}_{p}=\varepsilon^{2}$) thanks to the continuity equation and the divergence constraint $\nabla\cdot B=0$. As for the vorticity, taking curl in the momentum equation yields the evolution equation of vorticity
@@ 2.3 Reduction of pressure and design of energy functional
${\|(u,B)\|_{4}}$${\|\varepsilon^{2}\mathcal{T}^{2}(u,B)\|_{3}}$${\|\varepsilon^{4}\mathcal{T}^{4}(u,B)\|_{2}}$${\|\varepsilon^{6}\mathcal{T}^{6}(u,B)\|_{1}}$${\|\varepsilon^{8}\mathcal{T}^{8}(u,B)\|_{0}}$${\|\varepsilon^{2}\mathcal{T}p\|_{3}}$${\|\varepsilon^{4}\mathcal{T}^{3}p\|_{2}}$${\|\varepsilon^{6}\mathcal{T}^{5}p\|_{1}}$${\|\varepsilon^{8}\mathcal{T}^{7}p\|_{0}}$reduction of ${\text{reduction of }E_{4}}$reduction of ${\text{reduction of }E_{5}}$reduction of ${\text{reduction of }E_{6}}$reduction of ${\text{reduction of }E_{7}}$reduction of ${\text{reduction of }E_{8}}$closed  div  curl  div  curl  div  curl  div  curl  $\scriptstyle{\partial_{t}\text{ appears}}$  $\scriptstyle{\partial_{t}\text{ appears}}$  $\scriptstyle{\partial_{t}\text{ appears}}$  $\scriptstyle{\partial_{t}\text{ appears}}$  $\scriptstyle{\partial_{t}\text{ appears}}$      tangential      tangential      tangential      tangential  tangential (2.7)
@@
${E_{5}(t)}$${\|(u,B)\|_{4}}$${\|\partial_{t}(u,B)\|_{3}}$${\|\varepsilon\partial_{t}^{2}(u,B)\|_{2}}$${\|\varepsilon^{2}\partial_{t}^{3}(u,B)\|_{1}}$${\|\varepsilon^{3}\partial_{t}^{4}(u,B)\|_{0}}$${\|p\|_{4}}$${\|\varepsilon\partial_{t}p\|_{3}}$${\|\varepsilon^{2}\partial_{t}^{2}p\|_{2}}$${\|\varepsilon^{3}\partial_{t}^{3}p\|_{1}}$${\|\varepsilon^{4}\partial_{t}^{4}p\|_{0}}$closed  curl  curl  (d)  curl  (d)  curl  (d)  (c)  (c)  (c)  (c)  tangential (2.8)
@@ 3.1 $L^{2}$ estimate
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho|u|^{2}\,\mathrm{d}x=$ $~{}\int_{\Omega}(B\cdot\nabla)B\cdot u\,\mathrm{d}x-\int_{\Omega}u\cdot\nabla p\,\mathrm{d}x-\int_{\Omega}u\cdot\nabla(\frac{1}{2}|B|^{2})\,\mathrm{d}x$ (3.1)
$+\int_{\Omega}\frac{1}{2}\partial_{t}\rho|u|^{2}-\rho(u\cdot\nabla)u\cdot u\,\mathrm{d}x$
$=$ $-\int_{\Omega}(B\cdot\nabla)u\cdot B+\int_{\Omega}p(\nabla\cdot u)\,\mathrm{d}x+\frac{1}{2}\int_{\Omega}(\nabla\cdot u)|B|^{2}\,\mathrm{d}x$
$+\int_{\Omega}\underbrace{\frac{1}{2}\partial_{t}\rho|u|^{2}+\frac{1}{2}\nabla\cdot(\rho u)|u|^{2}}_{=0}\,\mathrm{d}x.$
Then, inserting the continuity equation and the evolution equation of $B$, we get
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho|u|^{2}\,\mathrm{d}x=$ $-\int_{\Omega}\partial_{t}B\cdot B-\int_{\Omega}(u\cdot\nabla)B\cdot B-\int_{\Omega}|B|^{2}(\nabla\cdot u)+\frac{1}{2}\int_{\Omega}(\nabla\cdot u)|B|^{2}\,\mathrm{d}x$ (3.2)
$-\int_{\Omega}\mathcal{F}_{p}p\partial_{t}p\,\mathrm{d}x-\int_{\Omega}\mathcal{F}_{p}p(u\cdot\nabla)p\,\mathrm{d}x$
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}|B|^{2}+\mathcal{F}_{p}|p|^{2}\,\mathrm{d}x+\frac{1}{2}\int_{\Omega}((\nabla\cdot u)+(\partial_{t}+u\cdot\nabla)\mathcal{F}_{p})|p|^{2}\,\mathrm{d}x.$
Similarly, we have $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho S^{2}=0$ thanks to $D_{t}S=0$, so we get the $L^{2}$ estimate:
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho|u|^{2}+|B|^{2}+\mathcal{F}_{p}|p|^{2}+\rho S^{2}\,\mathrm{d}x\leq P(E_{4}(t)).$ (3.3)
@@ 3.2.1 Reduction of pressure
Recall that $D_{t}=\partial_{t}+\bar{u}\cdot\overline{\nabla}+u_{3}\partial_{3}$ and $B\cdot\nabla=\bar{B}\cdot\overline{\nabla}+B_{3}\partial_{3}$ are both tangential derivatives thanks to $u_{3}=B_{3}=0$ on $\Sigma$. The above inequalities (3.5)-(3.6) show that $\nabla P$ is reduced to $\partial_{t}u$, ¯$\overline{\partial}{}u$, $\omega(x_{3})\partial_{3}u$, ¯$\overline{\partial}{}B$ and $\omega(x_{3})\partial_{3}B$ for some weight function $\omega(x_{3})$ vanishing on $\Sigma$. Thus, the momentum equation reduces a normal derivative of $P$ to a tangential derivative of $u,B$.
Also, one can reduce the Sobolev norm of $\nabla P$ in the same way. Let $\mathcal{T}=\partial_{t}$ or $\overline{\partial}{}$ or $\omega(x_{3})\partial_{3}$ and $D=\partial$ or $\partial_{t}$. The above estimate yields the control of $\|\mathcal{T}^{\alpha}D^{k}\nabla p\|_{0}$ after taking $\mathcal{T}^{\alpha}D^{k}$, $k+\langle\alpha\rangle\geq 1$, to (3.4). Specifically, at the leading order, $\|\mathcal{T}^{\alpha}D^{k}\nabla(p+|B|^{2}/2)\|_{0}$ is controlled by
@@ 3.2.2 Div-Curl analysis
$\|X\|_{s}^{2}\lesssim\|X\|_{0}^{2}+\|\nabla\cdot X\|_{s-1}^{2}+\|\nabla\times X\|_{s-1}^{2}+|X\cdot N|^{2}_{s-\frac{1}{2}}.$ (3.8)
@@
The $L^{2}$ norm of $u$ and $B$ has been controlled in (3.3), while the control of $\|\varepsilon^{(k-1)_{+}}\partial_{t}^{k}(u,B)\|_{0}~{}(1\leq k\leq 3)$ is parallel to the case of $k=0$ and will be postponed to later sections. The divergence part is reduced to the estimates of $p$ by using the continuity equation and (1.9)
$\|\varepsilon^{(k-1)_{+}}\nabla\cdot\partial_{t}^{k}u\|_{3-k}^{2}\lesssim\left\|\varepsilon^{(k-1)_{+}+2}\partial_{t}^{k}(D_{t}p)\right\|_{3-k}^{2},$ (3.11)
@@
It holds that
$\sum_{k=0}^{3}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}u\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}B\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}}\mathcal{F}_{p}^{\frac{1}{2}}B\times(\nabla\times\partial_{t}^{k}B)\right\|_{3-k}^{2}\right)\leq P(E_{4}(t),E_{5}(t)).$ (3.12)
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho\left|\partial^{3}\nabla\times u\right|^{2}+\left|\partial^{3}\nabla\times B\right|^{2}+\left|\mathcal{F}_{p}^{\frac{1}{2}}B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t),E_{5}(t)).$
@@
$\underbrace{+\int_{\Omega}(\partial^{3}\nabla\times B)\cdot\partial^{3}\nabla\times(\mathcal{F}_{p}BD_{t}p)\,\mathrm{d}x}_{:=\mathcal{K}}.$ (3.17)
Based on the concrete form of commutators, a straightforward product estimate for $\mathcal{I}_{j}$ $(1\leq j\leq 4)$ gives
@@
The most difficult term is $\mathcal{K}$ because the highest-order term has 5-th order derivative and thus cannot be controlled by $E_{4}(t)$ (we omit the terms generated by derivatives falling on $\mathcal{F}_{p}$ because they must be lower-order and have no loss of $\varepsilon$-weight thanks to (1.9) and $\partial_{t}S=O(1)$):
$\mathcal{F}_{p}\partial^{3}\nabla\times(BD_{t}p)=$ $~{}\mathcal{F}_{p}\partial^{3}((\nabla D_{t}p)\times B+D_{t}p(\nabla\times B))$
$=$ $-{\mathcal{F}_{p}}B\times(\partial^{3}\nabla D_{t}p)\underbrace{-{\mathcal{F}_{p}}[\partial^{3},B\times](\nabla D_{t}p)+\partial^{3}(\mathcal{F}_{p}D_{t}p(\nabla\times B))}_{\mathcal{R}_{2}},$
where straightforward calculation shows that $\|\mathcal{R}_{2}\|_{0}\leq P(E_{4}(t))$. We then further analyze the problematic term $-\mathcal{F}_{p}B\times(\partial^{3}\nabla D_{t}p)$. We invoke a simple vector identity (in 3D) $(B\cdot\nabla)B-\frac{1}{2}\nabla|B|^{2}=-B\times(\nabla\times B)$ to rewrite the momentum equation to be
@@
$-\mathcal{F}_{p}B\times(\partial^{3}\nabla D_{t}p)=-{\mathcal{F}_{p}}B\times(\partial^{3}D_{t}\nabla p)-\mathcal{F}_{p}B\times(\partial^{3}(\nabla u_{j}\partial_{j}p))$
$=$ $~{}\mathcal{F}_{p}B\times\partial^{3}D_{t}(\rho D_{t}u)+{\mathcal{F}_{p}}B\times\partial^{3}D_{t}(B\times(\nabla\times B))-\mathcal{F}_{p}B\times(\partial^{3}(\nabla u_{j}\partial_{j}p))$
$=$ $~{}\mathcal{F}_{p}\rho B\times(\partial^{3}D_{t}^{2}u)+\underbrace{{\mathcal{F}_{p}}B\times D_{t}(B\times\partial^{3}(\nabla\times B))}_{=\mathcal{K}_{1}}$
$\underbrace{+\mathcal{F}_{p}B\times([\partial^{3}D_{t},\rho]D_{t}u)+\mathcal{F}_{p}B\times\left([\partial^{3},D_{t}](\nabla\times B)+D_{t}\left([\partial^{3},B\times](\nabla\times B)\right)\right)-\mathcal{F}_{p}B\times(\partial^{3}(\nabla u_{j}\partial_{j}p))}_{\mathcal{R}_{3}},$ (3.19)
@@
$\int_{\Omega}\mathcal{F}_{p}(\partial^{3}\nabla\times B)\cdot\left(B\times D_{t}\left(B\times\partial^{3}(\nabla\times B)\right)\right)\,\mathrm{d}x.$ (3.20)
@@
$\int_{\Omega}\mathcal{F}_{p}(\partial^{3}\nabla\times B)\cdot\left(B\times D_{t}\left(B\times\partial^{3}(\nabla\times B)\right)\right)\,\mathrm{d}x=-\int_{\Omega}{\mathcal{F}_{p}}D_{t}\left(B\times(\partial^{3}\nabla\times B)\right)\cdot\left(B\times(\partial^{3}\nabla\times B)\right)\,\mathrm{d}x,$ (3.21)
@@
$-\int_{\Omega}{\mathcal{F}_{p}}D_{t}\left(B\times(\partial^{3}\nabla\times B)\right)\cdot\left(B\times(\partial^{3}\nabla\times B)\right)\,\mathrm{d}x$
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{\mathcal{F}_{p}}\left|B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x+\underbrace{\frac{1}{2}\int_{\Omega}(\nabla\cdot u+{D_{t}\mathcal{F}_{p}})\left|B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x}_{\mathcal{R}_{4}}.$ (3.22)
@@
$\mathcal{K}+\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{\mathcal{F}_{p}}\left|B\times(\partial^{3}\nabla\times B)\right|^{2}\,\mathrm{d}x\leq\|B\|_{4}\left(P(E_{4}(t))+\|\rho B\|_{\infty}\|\varepsilon^{2}D_{t}^{2}u\|_{3}\right)\leq P(E_{4}(t),E_{5}(t)),$ (3.23)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\|\nabla\times u\|_{3}^{2}+\|\nabla\times B\|_{3}^{2}+\left\|{\mathcal{F}_{p}}^{\frac{1}{2}}B\times(\nabla\times B)\right\|_{3}^{2}\right)\leq P(E_{4}(t),E_{5}(t)).$ (3.24)
@@
Similarly, we can prove the same conclusion for $\partial^{\alpha}\partial_{t}^{k}$ with $k+|\alpha|=3$ by replacing $\partial^{3}$ with $\varepsilon^{(k-1)_{+}}\partial^{\alpha}\partial_{t}^{k}$. Indeed, the highest order derivatives in the above commutators do not exceed 4-th order, and there is no loss of Mach number weight because none of the above steps creates negative power of Mach number. Besides, the problematic term can still be analyzed in the same manner. Throughout this section, we omit the terms generated by derivatives falling on $\mathcal{F}_{p}$.
$\varepsilon^{(k-1)_{+}+2}B\times(\partial^{\alpha}\partial_{t}^{k}\nabla D_{t}p)\stackrel{{\scriptstyle L}}{{=}}-\varepsilon^{(k-1)_{+}}\rho B\times\partial^{\alpha}\partial_{t}^{k}(\varepsilon^{2}D_{t}^{2}u)-\varepsilon^{(k-1)_{+}}{\mathcal{F}_{p}}B\times D_{t}\left(B\times(\partial^{\alpha}\partial_{t}^{k}(\nabla\times B))\right),$
where the $L^{2}$ norms of the omitted lower-order terms are controlled by $P(E_{4}(t))$. The contribution of the last term in the above equality is again the energy term $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}{\mathcal{F}_{p}}\varepsilon^{(k-1)_{+}}\left|B\times(\partial^{\alpha}\partial_{t}^{k}\nabla\times B)\right|^{2}\,\mathrm{d}x$. The proof of this follows in the same way if we replace $\partial^{3}$ above by $\varepsilon^{(k-1)_{+}}\partial^{\alpha}\partial_{t}^{k}$ (using the vector identity $(\mathbf{u}\times\mathbf{v})\cdot\mathbf{w}=-(\mathbf{u}\times\mathbf{w})\cdot\mathbf{v}$. Hence, we conclude the vorticity analysis for $E_{4}(t)$ with the following inequality
$\sum_{k=0}^{3}\frac{\mathrm{d}}{\mathrm{d}t}\left(\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}u\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}}\nabla\times\partial_{t}^{k}B\right\|_{3-k}^{2}+\left\|\varepsilon^{(k-1)_{+}}{\mathcal{F}_{p}^{\frac{1}{2}}}B\times\left(\partial_{t}^{k}\nabla\times B\right)\right\|_{3-k}^{2}\right)\leq P(E_{4}(t),E_{5}(t)).$ (3.25)
@@
[…]